I've been having problems with this particular problem:
A person won a state lottery prize of $20,000, from which 25% was deducted for taxes. The remainder was invested, partly for a 40% gain and a 10% loss. How much was each investment if there as a $2000 net investment gain?

I thought that you set up the problem as an equation x being the unknown amount but I got stumped on whether or not to use the pre or post taxed amount.

Answer by stanbon(75887)

A person won a state lottery prize of $20,000, from which 25% was deducted for taxes. The remainder was invested, partly for a 40% gain and a 10% loss. How much was each investment if there as a $2000 net investment gain?
----------------------------------
From the lottery he received 0.75*20000 = 15000
---------------------------------------------------
Equation:
interest + interest = interest
0.40x - 0.10(15000-x) = 2000
Multiply thru by 100 to get:
40x - 150000 + 10x = 200000
50x = 350000
x = $7000 (amt. invested at 40%)
15000-7000 = $8000 (amt. invested at 10%)
==============================================
Cheers,
Stan H.